Suppose 8 pints of a 13% alcohol solution is mixed with 2 pints of a 60% alcohol solution. What is the concentration of alcohol in the new 10 pint mixture?
Found 2 solutions by checkley79, richwmiller:
Answer by checkley79(3341)

.13*8+.60*2=10x
1.04+1.2=10x
2.24=10x
x=2.24/10
x=.224% mixture.

Answer by richwmiller(17219)

.13*8+.60*2=10x
1.04+1.2=10x
2.24=10x
x=2.24/10
x=.224
or
22.4%
not .224%